OpenStudy (anonymous):

The answer to (2/3) +(1/3) (4/5) = 4/5 A. True B. False

OpenStudy (anonymous):

2/3 + 1/3 = 3/3 which is a whole number 1. 4/5 = 4/5 which is not a whole number so it's FALSE.